OpenStudy (anonymous):

Mel Company has a net income, before taxes, of $95,000. The treasurer of the company estimates 45 percent of net income will have to be paid for federal and state taxes. The tax for both federal and state is

OpenStudy (anonymous):

45% = taxes and stuff 45 % of 95,000 = 45 /100 * 95,000 = 45 * 950 = 42,750